### v1

# Choosing alpha post hoc: the danger of multiple standard significance thresholds

Abstract
A fundamental assumption of classical hypothesis testing is that the significance threshold $\alpha$ is chosen independently from the data. The validity of confidence intervals likewise relies on choosing $\alpha$ beforehand. We point out that the independence of $\alpha$ is guaranteed in practice because, in most fields, there exists one standard $\alpha$ that everyone uses – so that $\alpha$ is automatically independent of everything. However, there have been recent calls to decrease $\alpha$ from $0.05$ to $0.005$. We note that this may lead to multiple accepted standard thresholds within one scientific field. For example, different journals may require different significance thresholds. As a consequence, some researchers may be tempted to conveniently choose their $\alpha$ based on their p-value. We use examples to illustrate that this severely invalidates hypothesis tests, and mention some potential solutions.
1 Introduction
1.1 The call to decrease significance thresholds
Recently, there has been an intense debate in many scientific fields on whether the significance threshold for declaring scientific discoveries should be changed from $0.05$ to $0.005$ (Borenstein, 2001; Ioannidis, 2005; Banerjee et al., 2009; Mudge et al., 2012; Martin and Bridgmon, 2012; Kim and Ji, 2015; Ioannidis, 2016; Szucs and Ioannidis, 2017; Benjamin et al., 2018; Ioannidis, 2018; Held, 2019; Di Leo and Sardanelli, 2020; Khan et al., 2020; Aguinis et al., 2021; Schnog et al., 2021; Thakur and Jha, 2022; Maier and Lakens, 2022; Fitzpatrick et al., 2024). In most statistics textbooks, the “significance threshold” or “significance level” is denoted by $\alpha$ and is assumed to be prespecified, i.e., data-independent (Van der Vaart, 1998; Lehmann and Romano, 2022; Hansen, 2022). In particular, $\alpha$ is assumed independent of the p-value, since otherwise $\alpha$ would depend on the data (Hubbard and Bayarri, 2003). One important paper that stimulated the mentioned debate, bears the title “Redefine statistical significance” (Benjamin et al., 2018). In this article, 72 influential scientists propose changing the “threshold for statistical significance for claims of new discoveries” to $0.005$.
As a consequence of this proposal, scientists may find themselves in a field where there is not one standard threshold, but multiple: some groups of scientists may support $\alpha=0.005$ and some $\alpha=0.05$. This can also translate into the implicit or explicit preferences of journals, with some journals preferring $0.005$ and others $0.05$.
1.2 The statistical danger of differing thresholds
The message of this paper is that if within one field, there are multiple standard significance thresholds, then researchers may be tempted to first analyze their data and then choose the $\alpha$ that suits them. The problem can best be indicated with an example.
Suppose a researcher is interested in a certain null hypothesis and plans to compute a p-value for it. In her field there are two standard significance thresholds, $0.005$ and $0.05$, simply because some researchers or journal editors in her field advocate 0.005 and others advocate 0.05. After she has computed the p-value, she wants to share it in some form, e.g. by presenting it at a conference or submitting it to a journal. Then the choice of where to present or where to submit her results will potentially depend on the p-value that she finds. For example, if she finds $p=0.02$, then she may be more likely to submit to a journal that accepts $\alpha=0.05$ than to a journal that prefers $\alpha=0.005$. On the other hand, if she finds an extremely small p-value, then she may prefer a journal with $\alpha=0.005$, since then her finding is significant at a more stringent threshold. Consequently, it would be tempting for the researcher not to fix her $\alpha$ before running her experiment, but to let her $\alpha$ depend on the journal or conference, which she chooses after obtaining her p-value.
The statistical issue that arises, is that the researcher’s $\alpha$ is no longer independent of her p-value. Consequently, the test loses its usual decision-theoretic interpretation, i.e., there is no guarantee anymore that the type I error probability is below $\alpha$ (Hubbard and Bayarri, 2003; Grünwald, 2024; Koning, 2024a). Note that this issue is about more than publication bias. It is about the validity of an individual test, from the perspective of the individual researcher: when she lets $\alpha$ depend on the data, she cannot validly interpret her own statistical test anymore.
In theory, a solution would be that the researcher fixes her significance threshold beforehand and never changes it. However, one can easily imagine a researcher who finds $p=0.003$ and then claims her significance level is $\alpha=0.005$, even though she never fixed $\alpha$ beforehand.
1.3 A related discussion: abandoning significance thresholds
In parallel to the discussion on what the threshold should be, there has been a discussion on whether significance thresholds should be used at all (Hubbard and Lindsay, 2008; Wasserstein and Lazar, 2016; McShane et al., 2019; Amrhein et al., 2019). Indeed, a downside of significance testing in practice is that it may overemphasize a specific p-value threshold rather than interpreting p-values continuously, considering effect sizes, and using other scientifically relevant information (Borenstein, 1994; Greenland et al., 2016; Thompson, 2007; Betensky, 2019; Scheel et al., 2021; Imbens, 2021; Muff et al., 2022). For example, the famous ASA statement on p-values (Wasserstein and Lazar, 2016) notes that “Scientific conclusions and business or policy decisions should not be based only on whether a p-value passes a specific threshold.” Note that the ASA statement is not against p-values, but against too much focus on p-value thresholds.
Many statisticians agree with the ASA statement’s point, but nevertheless we see that in the recent statistical literature, significance testing is still prominent (Lehmann and Romano, 2022; Hansen, 2022). One reason is that from a mathematical point of view, there is often essentially no difference between a formula for a significance test, a formula for a p-value and a formula for a confidence interval. For example, confidence intervals can often be directly derived by ‘inverting’ a test. Another reason is that some statisticians still value type I error control, which is only possible if some $\alpha$ is chosen. Hence, although many statisticians largely agree with the ASA statement, they continue to develop hypothesis tests.
Beyond theoretical statistics, in science in general we see that significance testing has not disappeared (Hubbard, 2019; Mayo and Hand, 2022; Thakur and Jha, 2022; Maier and Lakens, 2022; Hansen, 2022; Fitzpatrick et al., 2024). Moreover, if one abandons significance tests, then one is still likely to compute confidence intervals, which also require choosing $\alpha$. In that case, choosing $\alpha$ post hoc causes issues similar to those discussed in this paper.
1.4 This paper
Throughout the paper, we stay within the classical Neyman-Pearson theory of hypothesis testing. We note that one of the main reasons why classical Neyman-Pearson hypothesis testing works in practice, is that there exists a single standard level $\alpha$ in most fields. This automatically avoids that $\alpha$ is chosen based on the data. Although it is known that having $\alpha$ depend on the p-value is invalid (Hubbard and Bayarri, 2003), this paper seems to be the first to point out this will likely occur when within a field there are multiple standard thresholds.
Further, we discuss in detail what can go wrong in terms of type I error control. We evaluate what the type I error rate is, conditional on a certain $\alpha$ being chosen based on the p-value. Apart from discussing the situation where the researcher chooses between two significance levels, we will also consider the most extreme scenario where there are many choices for $\alpha$.
Moreover, for a test with a data-dependent level $\alpha$, we consider a measure of evidence for how problematic this is on average. If this measure is large, this means the conditional type I error rate of the test is much too large in an average sense. In the examples that we provide in Sections 3 and 4, the interpretation of the hypothesis test is ruined due to the post hoc selection of $\alpha$.
Throughout, we focus on the case where one hypothesis is tested. In case there are multiple hypotheses and a multiple testing correction is used, similar issues arise when $\alpha$ is chosen post hoc. This is discussed in the Supplementary Material. In Section 5 and the Discussion, we briefly point out some potential solutions to the discussed issue.
2 Notation and concepts
2.1 Tests with data-dependent $\alpha$
Consider any hypothesis $H_{0}$ and a corresponding hypothesis test $\phi_{\alpha}$. The subscript is used because the test depends on the chosen significance level $\alpha>0$. The test $\phi_{\alpha}$ is a function from the sample space to the set $\{0,1\}$, where $1$ indicates that $H_{0}$ is rejected and 00 indicates that $H_{0}$ is not rejected. We will indicate the data with $X$, which takes values in the sample space. We view the test $\phi_{\alpha}=\phi_{\alpha}(X)$ as a function of $X$, i.e., given $\alpha$, the outcome of the test $\phi_{\alpha}$ is entirely determined by the data $X$.
In classical hypothesis testing, the level $\alpha$ is always assumed to be selected independently from the data. For such an independently selected $\alpha$, it is typical to use a test $\phi_{\alpha}$ with the guarantee that it does not reject the hypothesis too often when it is true, namely with probability at most $\alpha$. We assume throughout that if $\alpha$ is chosen independently from $X$, then the tests we consider are exactly valid, in the sense that $\mathbb{P}(\phi_{\alpha}=1)=\alpha$ if $H_{0}$ is true.
As indicated in the introduction, there are scenarios where $\alpha$ would depend on $X$ – for example if $\alpha$ depends on the p-value. Then $\alpha$ is a function of $X$, although we will often write $\alpha$ instead on $\alpha(X)$ for short. If $\alpha$ is data-dependent, then conditional on $\alpha$ taking a certain value, we possibly lose the guarantee that $\mathbb{P}_{H_{0}}(\phi_{\alpha}=1)\leq\alpha$. For example, suppose $\alpha$ takes two possible values, e.g. 0.005 and 0.05, depending on the data $X$. Then, depending on how $\alpha$ exactly depends on $X$, it can be the case that $\mathbb{P}({\phi_{\alpha}=1|\alpha=0.005})$ is not at most $0.005$, but larger than $0.005$. This will be illustrated in Section 3.
2.2 Measuring the discrepancy between $\alpha$ and the type I error rate
Let $A\subseteq[0,1]$ denote the set of possible values that $\alpha(X)$ can take. Throughout we use ‘$a$’ to denote a value from $A$, independent of the data. We can then ask whether conditional on the data-dependent $\alpha$ taking the value $a$, the type I error rate is below $\alpha$, i.e. whether $\mathbb{P}_{H_{0}}(\phi_{\alpha}=1|\alpha=a)\leq a$. A measure of the discrepancy between $a$ and $\mathbb{P}_{H_{0}}(\phi_{\alpha}=1|\alpha=a)$ is the difference
$d_{a}:=\mathbb{P}\big{(}\phi_{\alpha}=1|\alpha=a\big{)}-a.$ (1)
Ideally, we would have $d_{a}=0$ for all $a\in A$, but this is often not the case for many $a\in A$, if $\alpha$ depends on the data. This is well known (Hubbard and Bayarri, 2003) and illustrated in Sections 3 and 4. Now suppose that for some $a\in A$, the difference is $d_{a}=0.01$. Do we consider this difference small or large? To a large extent, this depends on what $a$ is. For example, if $a=0.05$, then having $d_{a}=0.01$ is not highly problematic, since $\mathbb{P}(\phi_{\alpha}=1|\alpha=a)=0.06$ is then not much larger than $a$, in relative terms. However, if $a=0.005$, then $d_{a}=0.01$ implies that the true error rate $\mathbb{P}_{H_{0}}(\phi_{\alpha}=1|\alpha=a)=0.015$ is three times larger than $a$. In that case, $d_{a}=0.01$ is a huge difference. Thus, $d_{a}$ is arguably not the best measure of how liberal the test is. Instead it can be more meaningful to look at the relative discrepancy
$r_{a}:=\mathbb{P}_{H_{0}}\big{(}\phi_{\alpha}=1|\alpha=a\big{)}/a,$ (2)
where we remind the reader that $\alpha=\alpha(X)$ depends on the data. We will call $r_{a}$ the discrepancy ratio conditional on $\alpha=a$. We would like this ratio to be at most 1. If $r_{a}<1$, then the test is conservative conditional on $\alpha=a$, and if $r_{a}>1$ the test is liberal conditional on $\alpha=a$. Note that for a given number $a$, $r_{a}$ is a fixed number. In contrast, since $\alpha=\alpha(X)$ depends on the data, $r_{\alpha}$ is data-dependent.
2.3 Expected discrepancy
If the discrepancy ratio $r_{a}$ is below 1 for every $a\in A$, then in particular the expected value of the discrepancy ratio will be below 1, i.e.,
$\mathbb{E}r_{\alpha}=\mathbb{E}\big{\{}\mathbb{P}(\phi_{\alpha}=1|\alpha)/\alpha\big{\}}$ (3)
will be at most 1. Note that this is the expected value of a conditional probability divided by $\alpha$. Likewise, if (3) is larger than 1, it follows that there is at least one $a\in A$ for which (2) is larger than 1.
Note that in general, if $\mathbb{E}r_{\alpha}$ is much larger than $1$ under $H_{0}$, it follows that either there is at least a small probability that $r_{\alpha}$ is much larger than $1$, or there is a large probability that $r_{\alpha}$ is at least somewhat larger than 1. Both these possibilities are problematic for researchers. Thus, $\mathbb{E}r_{\alpha}$ is a summary measure of how wrong things go on average. If $\mathbb{E}r_{\alpha}\gg 1$, this indicates that the test procedure is seriously flawed, in an average sense.
It is useful to note that $\mathbb{E}r_{\alpha}$ can be rewritten in the following ways:
$\mathbb{E}r_{\alpha}=\mathbb{E}\big{\{}\mathbb{P}(\phi_{\alpha}=1|\alpha)/\alpha\big{\}}=\mathbb{E}\big{\{}\mathbb{E}\big{(}\phi_{\alpha}|\alpha\big{)}/\alpha\big{\}}=$
$\mathbb{E}\big{\{}\mathbb{E}\big{(}\phi_{\alpha}/\alpha|\alpha\big{)}\big{\}}=\mathbb{E}\big{(}\phi_{\alpha}/\alpha\big{)}.$ (4)
The equalities (4) tell us that $\mathbb{E}r_{\alpha}$ can be rewritten as the expectation of a conditional expectation, which can in turn be rewritten as a simple expectation.
3 Example 1: two possible significance levels $\alpha$
Consider a researcher who tests one hypothesis. In her field there are two standard significance thresholds, $a_{1}$ and $a_{2}$, where $10<a_{1}<a_{2}<$ 1. As explained in the Introduction, there are scenarios where her $\alpha$ would depend on the p-value. In particular, if the researcher obtains a p-value $p$ with $a_{1}<p\leq a_{2}$, she may be more likely to take $\alpha=a_{2}$ than when $p\leq a_{1}$, in which case she could be more likely to take $\alpha=a_{2}$. As a simple behavioural model, suppose that the researcher takes $\alpha=a_{1}$ if $p\leq a_{1}$ and takes $\alpha=a_{2}$ if $a_{1}<p\leq 1$.
We will use the notation from Section 2.1. Since the threshold depends on the p-value, it depends on the data. Thus, $\alpha(X)$ depends on $X$, although we supress ‘$(X)$’ for brevity. We have $\phi_{\alpha}=\phi_{\alpha(X)}=1$ if and only if $p\leq\alpha$. If $\alpha$ would be fixed, then the test $\phi_{\alpha}$ would be valid in the sense that $\mathbb{P}(\phi_{\alpha}=1)\leq\alpha$. However, because $\alpha$ depends on the p-value, the interpretation of the test is ruined, as we will see.
3.1 No type I error control conditional on $\alpha$ or marginally
We investigate the type I error probability conditional on a certain value for $\alpha$ being chosen. What happens conditional on $\alpha=a_{1}$ can be seen immediately: if $\alpha=a_{1}$, then this means that $p$ was apparently at most $a_{1}$. This means that $H_{0}$ is rejected. Thus, under $H_{0}$, conditional on $\alpha=a_{1}$, the type I error probability is not $a_{1}$, but 1. Further, the type I error probability is larger than $a_{1}$ marginally. Thus, the familiar interpretation of the hypotheses test is ruined. This means that the researcher cannot validly interpret her own statistical test anymore, in the usual decision-theoretic way.
The following proposition states what happens to the type I error probability conditional on $\alpha=a_{2}$. The proof is in the appendix.
Proposition 1.
For the above procedure that takes $\alpha=a_{1}$ if $p\leq a_{1}$ and $\alpha=a_{2}$ otherwise, under $H_{0}$ we have
$\mathbb{P}(\phi_{\alpha}=1|\alpha=a_{2})=\big{(}(1-a_{1}/a_{2})/(1-a_{1})\big{)}a_{2}<a_{2}.$
Thus, conditional on the data-dependent $\alpha$ being $a_{2}$, the type I error rate does not exceed $a_{2}$, although the test is conservative since $\mathbb{P}(\phi_{\alpha}=1)$ is strictly smaller than $a_{2}$. For example, for $a_{1}=0.005$ and $a_{2}=0.05$, we have $\mathbb{P}(\phi_{\alpha}=1|\alpha=a_{2})=a_{2}((1-0.005/0.05)/(1-0.005))\approx a_{2}\cdot 0.9045$. Thus, in the notation of Section 2.2, $r_{a_{2}}\approx 0.9045$.
In summary, conditional on $\alpha=a_{1}$ the type I error probability is too large and conditional on $\alpha=a_{2}$ the type I error probability is on the conservative side. Put differently, the discrepancy ratio is $r_{\alpha}>1$ if $\alpha=a_{1}$ and $r_{\alpha}<1$ if $\alpha=a_{2}$.
3.2 No type I error control in terms of expected discrepancy
Next, we can ask whether perhaps we control the type I error probability in expectation, i.e., whether $\mathbb{E}r_{\alpha}\leq 1$ under $H_{0}$. This would mean that at least in an average sense, we have type I error control. The following proposition implies that this does not hold.
Proposition 2.
For the procedure that takes $\alpha=a_{1}$ if $p\leq a_{1}$ and $a_{2}$ otherwise, under $H_{0}$ we have
$\mathbb{E}r_{\alpha}=1+(a_{2}-a_{1})/a_{2}>1.$
For example, if $a_{1}=0.005$ and $a_{2}=0.05$, then $\mathbb{E}_{H_{0}}(r_{\alpha})=1.9$. Since $\mathbb{E}_{H_{0}}(r_{\alpha})$ is not bounded by 1, it follows that $r_{\alpha}$ is larger than 1 conditional on at least one possible value of $\alpha$. We already saw that this is the case for $\alpha=a_{1}$. We conclude that both conditionally and on average, the researcher’s test does not provide type I error control.
4 Example 2: Infinitely many significance levels $\alpha$
Now suppose there are not two thresholds to choose from but many. To take this to the extreme, suppose each $0<\alpha\leq C$ can be chosen. Here $0<C\leq 1$ may for example be $0.1$. It is then easily imaginable that the author is biased towards choosing an $\alpha$ that is at least as large as her p-value $p$, when possible. To fix ideas, suppose the researcher in fact selects $\alpha=p$, unless $p>C$, in which case she selects $\alpha=C$. As usual, we denote the test by $\phi_{\alpha}$, where $\alpha$ depends on $p$.
In case $p\leq C$, we always have $p\leq\alpha$, so that $H_{0}$ is always rejected. Likewise, if $\alpha<C$, $H_{0}$ is always rejected. Thus, under $H_{0}$, conditional on $\alpha<C$, the type I error rate is not below $\alpha$. Specifically, for every $a<C$, we have $r_{a}=\mathbb{E}(\phi_{\alpha}|\alpha=a)/a=1/a$. Thus, the smaller $a$ is, the larger $r_{a}$ is, i.e. the larger discrepancy ratio is conditional on $\alpha=a$.
Next, we may ask whether we have type I error control not conditionally, but at least on average, in the sense that $\mathbb{E}(r_{\alpha})\leq 1$ under $H_{0}$. Since $p$ is uniform on $[0,1]$ and $p\leq\alpha$ if and only if $p\leq C$, we have
$\mathbb{E}(r_{\alpha})=\mathbb{E}(\phi_{\alpha}/\alpha)=$
$\mathbb{P}(p>\alpha)\cdot 0+\mathbb{P}(p\leq\alpha)\mathbb{E}(\phi_{\alpha}/\alpha|p\leq\alpha)=$
$\mathbb{P}(p\leq\alpha)\mathbb{E}(\phi_{\alpha}/\alpha|\alpha=p\leq C)=$
$\mathbb{P}(p\leq\alpha)\mathbb{E}(1/p|p\leq C)=$
[…]
We see that the expected discrepancy ratio $\mathbb{E}(r_{\alpha})$ is not only larger than 1, but in fact infinity. Thus, in expectation, the test is completely out of control.
5 Connection to e-values
A partial solution to the problem of data-dependent $\alpha$ are so-called e-values. An e-value is a measure of evidence that has been recently proposed as an alternative to the p-value (Howard et al., 2021; Shafer, 2021; Vovk and Wang, 2021; Grünwald et al., 2024; Ramdas et al., 2023; Koning, 2024b). An e-value is typically defined as a non-negative random variable $e$, whose expectation is bounded by 1 under the hypothesis $H_{0}$:
$\mathbb{E}\,e\leq 1.$ (5)
In the context of a simple null and alternative hypothesis that only contain a single distribution, a prominent example of an $e$-value is the likelihood ratio between these distributions (Grünwald et al., 2024; Larsson et al., 2024).
The reciprocal $p^{*}=1/e$ of an e-value gives rise to a conservative p-value. Indeed, every such $p$-value is valid:
$\mathbb{P}(p^{*}\leq\alpha)=\mathbb{P}(e\geq 1/\alpha)\leq\alpha\,\mathbb{E}\,e\leq\alpha,$
for all $\alpha>0$, where the first inequality follows from Markov’s inequality and the second inequality from the definition of the e-value.
Recently, Koning (2024a) has shown that such a p-value is not just some conservative p-value, but has deep links to the data-dependent selection of the level and the expected relative discrepancy (see also Grünwald, 2024). In particular, an e-value-based p-value $p^{*}$ leads to decisions that have an expected discrepancy ratio below 1 for every data-dependent choice of the level $\alpha$. That is, for every possibly data-dependent level $\alpha$,
$\mathbb{E}\,r_{\alpha}=\mathbb{E}\left[\frac{\mathbb{P}(p^{*}\leq\alpha)}{\alpha}\right]\leq 1.$ (6)
A consequence is that a test $\phi_{\alpha}=\mathbb{I}\{p^{*}\leq\alpha\}$ based such a special p-value $p^{*}$ is valid for any data-dependent $\alpha$, in the sense that the expected relative discrepancy remains below 1. This is in stark contrast to traditional p-values, for which this error can be $\infty$, as discussed in Section 4. The connection is that a traditional $p$-value only satisfies (6) for every data-independent level.
While e-values are useful when $\alpha$ depends on the data, they should not be used when $\alpha$ is already fixed in advance. Indeed, in that case there is a considerable power loss, compared to using a traditional p-value.
Discussion
The fundamental point of this paper is that significance testing works in practice because researchers all by default use the same $\alpha$. This ensures that $\alpha$ is independent of the data. If there are multiple standards for $\alpha$ in a field and a researcher chooses her $\alpha$ post hoc, then that spells the end of the validity of her tests. Her p-values may be perfectly valid and they will provide a pragmatic measure of the amount of proof against $H_{0}$ (Efron, 2010, p.31), but she will not have a test with the usual interpretation in terms of type I error control. Likewise, confidence intervals are no longer valid if $\alpha$ is chosen post hoc. As statisticians we should give this some careful thought. One solution would be that researchers fix $\alpha$ in advance and never change it, but they would not necessarily have enough incentives or awareness to do this.
When a community of researchers within a scientific field starts adopting a lower $\alpha$, say $\alpha=0.005$, there are two possible scenarios: 1. Everyone in that field adopts $\alpha=0.005$ at the same time; 2. some stick to $\alpha=0.05$ while others adopt $\alpha=0.005$. The problem described in this paper occurs in scenario 2.
Existing papers that advocate decreasing $\alpha$ (Johnson et al., 2017; Szucs and Ioannidis, 2017; Benjamin et al., 2018; Ioannidis, 2018; Held, 2019; Schnog et al., 2021; Fitzpatrick et al., 2024), do not discuss the following question: would they already consider it an improvement if some people or journals in a field decrease $\alpha$, or do they consider it essential that the field as a whole uses the same $\alpha$? The present paper illustrates that these distinctions matter.
Suppose that in a field the scenario occurs where some stick to $\alpha=0.05$ while others adopt $\alpha=0.005$ – thus providing an incentive for letting $\alpha$ depend on the data. One solution would be for journals to require pre-registration of the $\alpha$ before researchers conduct an experiment. Another partial solution would be to use e-values instead of p-values, as discussed in Section 5. Another way out is that a researcher uses a Bayesian approach, which does not require fixing an $\alpha$ and nevertheless tells us the probabilty that $H_{0}$ is false – which frequentist statistics never does (Kelter, 2020; Lakens et al., 2020; van de Schoot et al., 2021). A potential downside of Bayesian inference is that one needs to specify a prior distribution, which often depends on subjective considerations.
Appendix A Proofs
A.1 Proof of Proposition 1
Proof.
Since $\alpha=a_{2}\Leftrightarrow p>a_{1}$ and $p\leq a_{2}\Leftrightarrow\phi_{\alpha}=1$, under $H_{0}$ we have
$\mathbb{P}(\phi_{\alpha}=1|\alpha=a_{2})=$
$\mathbb{P}\big{\{}p\leq a_{2}|p>a_{1}\big{\}}=$
$\mathbb{P}\big{\{}a_{1}<p\leq a_{2}|p>a_{1}\big{\}}=$
$(a_{2}-a_{1})/(1-a_{1})=$
$a_{2}(1-a_{1}/a_{2})/(1-a_{1})<$
$a_{2}(1-a_{1})/(1-a_{1})=a_{2}.$
A.2 Proof of Proposition 2
Proof.
By (4), under $H_{0}$ we have
$r_{\alpha}=\mathbb{E}(\phi_{\alpha}/\alpha)=$
$\mathbb{P}(p\leq a_{1})\mathbb{E}(\phi_{\alpha}/\alpha|p\leq a_{1})+$
$\mathbb{P}(a_{1}<p\leq a_{2})\mathbb{E}(\phi_{\alpha}/\alpha|a_{1}<p\leq a_{2})=$
$a_{1}(1/a_{1})+(a_{2}-a_{1})(1/a_{2})=$
$1+(a_{2}-a_{1})/a_{2}.$
Appendix B Multiple hypotheses
Here we discuss scenarios where $\alpha$ is chosen post hoc and multiple hypotheses are tested. Consider the situation where a particular researcher tests multiple hypotheses $H_{1},...,H_{m}$ and computes corresponding p-values $p_{1},...,p_{m}$. To avoid a flood of false positive findings, it is common in this setting to apply a multiple testing corrrection. Here we will consider control of the familywise error rate (FWER) by use of the Bonferroni method.
The Bonferroni method takes the raw, i.e. unadjusted, p-values $p_{1},...,p_{m}$ as input and outputs adjusted p-values, which are simply $p_{1}^{b}=\min\{p_{1}\cdot m,1\},...,p_{m}^{b}=\min\{p_{m}\cdot m,1\}$. The method then rejects the hypotheses with adjusted p-values at most $\alpha$. Of course, an equivalent way to see which hypotheses are rejected is to simply check which of the raw p-values are at most $\alpha/m$.
The number of false positives, i.e. the number of true hypotheses that are rejected, if often denoted by $V$. The FWER is defined as $\mathbb{P}(V>0)$ and controlling the FWER means ensuring that $\mathbb{P}(V>0)\leq\alpha$. The Bonferroni method controls the FWER regardless of the dependence structure of the $m$ p-values and regardless of how many hypotheses are in fact false.
It should furter be remarked that the Bonferroni procedure can be made slightly more powerful through a stepwise procedure by Holm (Holm, 1979; Goeman and Solari, 2014). This method checks which hypotheses are rejected by Bonferroni and removes these from consideration. Then Bonferroni is applied to the hypotheses that are left, which can lead to additional rejections and this is repeated until a step is reached where there are no additional rejections. Here we will simply consider Bonferroni without the Holm improvement, since if Bonferroni does not control the FWER then by implication neither does Holm.
While there are many scenarios with data-dependent $\alpha$ where Bonferroni fails miserably, here we will focus on two scenarios, which bear slight resemblance to the scenarios considered in §2. As is usual, we will assume that the raw p-values corresponding to the true hypotheses are uniformly distributed on $[0,1]$. Further, for the sake of the below examples, we assume that $p_{1},...,p_{m}$ are independent of each other. Note that $FWER\leq\alpha$ is equivalent to $\mathbb{E}(\mathbbm{1}_{E})\leq\alpha$, where $E$ is defined to be the event that $V>0$. Further note that $\mathbb{E}(\mathbbm{1}_{E})\leq\alpha$ is equivalent to $\mathbb{E}(\mathbbm{1}_{E}/\alpha)\leq 1$, which we use below.
B.1 Example 1: two possible significance levels $\alpha$
As in the main paper, suppose the researcher chooses from two thresholds $10<a_{1}<a_{2}<$ 1. Let $p_{(1)}=\min\{p_{i}:1\leq i\leq m\}$ denote the smallest raw p-value and let $p_{(1)}^{b}=\min\{mp_{(1)},1\}$ be the smallest Bonferroni-adjusted p-value. Similar to before, one can image that a researcher is more likely to use the threshold $a_{1}$ if there is at least one rejection at $\alpha=a_{1}$, i.e. if $p_{(1)}^{b}\leq a_{1}$. To simplify, suppose she chooses $\alpha=a_{1}$ if $p_{(1)}^{b}\leq a_{1}$ and $\alpha=a_{2}$ otherwise, i.e. if $a_{1}<p_{(1)}^{b}\leq 1$. Like before, $\alpha$ depends on the p-value and thus the data, but we write $\alpha$ instead of $\alpha(X)$ for brevity.
B.1.1 No type I error control conditional on $\alpha$
Suppose all $m$ hypotheses are true and recall that $E=\{V>0\}$ is the event that there is at least one false positive finding. Then if $\alpha=a_{1}$, we have $p_{(1)}^{b}\leq a_{1}$. Thus $\mathbb{P}(E|\alpha=a_{1})$ is 1, i.e., the FWER is not controlled at all conditional on $\alpha=a_{1}$. The following proposition states that conditional on $\alpha=a_{2}$ it is the other way around, i.e. the test is conservative: $\mathbb{P}(E|\alpha=a_{2})<a_{2}$.
Proposition 3.
Suppose $m\geq 2$ and all hypotheses $H_{1},...,H_{m}$ are true. For the above procedure that takes $\alpha=a_{1}$ if $p\leq a_{1}$ and $\alpha=a_{2}$ otherwise, we have
$\mathbb{P}(E|\alpha=a_{2})=1-\Big{\{}(1-a_{2}/m)^{m}\big{/}(1-a_{1}/m)^{m}\Big{\}}<a_{2}.$
Proof.
Part 1. We have
$\mathbb{P}(E|\alpha=a_{2})=\mathbb{P}(p_{(1)}^{b}\leq a_{2}|p_{(1)}^{b}>a_{1})=\mathbb{P}(a_{1}<p_{(1)}^{b}\leq a_{2}|p_{(1)}^{b}>a_{1})=$
$\mathbb{P}(a_{1}<mp_{(1)}\leq a_{2}|mp_{(1)}>a_{1})=\mathbb{P}(a_{1}/m<p_{(1)}\leq a_{2}/m|p_{(1)}>a_{1}/m).$ (7)
If $m$ variables are independent and have a standard uniform distribution, the minimum of the $m$ variables has a $Beta(1,m)$ distribution (Hemerik et al., 2019), which has density function $f(x)=\{1/B(1,m)\}(1-x)^{m-1}$, where $B(1,m)$ is the normalization constant
$\int_{0}^{1}(1-x)^{m-1}dx=[-(1-x)^{m}/m]\Big{|}_{0}^{1}=0--1/m=1/m.$
Thus, $p_{(1)}$ has a $Beta(1,m)$ distribution, so that (7) equals
[…]
$\Big{\{}-(1-x)^{m}/m\Big{|}_{a_{1}/m}^{a_{2}/m}\Big{\}}\Big{/}\Big{\{}-(1-x)^{m}/m\Big{|}_{a_{1}/m}^{1}\Big{\}}=$
$\Big{\{}-(1-a_{2}/m)^{m}--(1-a_{1}/m)^{m}\Big{\}}\Big{/}\Big{\{}0--(1-a_{1}/m)^{m}\Big{\}}=$
$\Big{\{}(1-a_{1}/m)^{m}-(1-a_{2}/m)^{m}\Big{\}}\Big{/}\Big{\{}(1-a_{1}/m)^{m}\Big{\}}=$
$1-\Big{\{}(1-a_{2}/m)^{m}\big{/}(1-a_{1}/m)^{m}\Big{\}},$
as was to be shown.
Part 2. We now need to show that the above is smaller than $a_{2}$. This is equivalent to showing that
$(1-a_{2}/m)^{m}/(1-a_{1}/m)^{m}<1-a_{2},$
which is equivalent to
$(m-a_{2})^{m}(1-a_{2})^{-1}(m-a_{1})^{-m}<1.$
The left side is $1$ if $a_{2}=a_{1}$ and we need to show that it is strictly larger than 1 if $a_{2}>a_{1}$. We will do this by showing that the derivative of the function
$g(a_{2})=(m-a_{2})^{m}(1-a_{2})^{-1}(m-a_{1})^{-m}$
is strictly positive on $(0,1)$. We use the product rule and the chain rule to compute the derivative:
$\frac{d}{da_{2}}g(a_{2})=\Big{\{}(m-a_{2})^{m}(--(1-a_{2})^{-2})-m(m-a_{2})^{m-1}(1-a_{2})^{-1}\Big{\}}(m-a_{1})^{-m}.$
Thus, to show that $\frac{d}{da_{2}}g(a_{2})>0$, it suffices to show that
$(m-a_{2})^{m}(1-a_{2})^{-2}>m(m-a_{2})^{m-1}(1-a_{2})^{-1},$
i.e., that
$(m-a_{2})(1-a_{2})^{-1}>m.$
This is true since
$(1-a_{2}/m)(1-a_{2})^{-1}>1,$
so that we are done.
So far, we have seen that $\mathbb{P}(E|\alpha=a_{2})<a_{2}$ and $\mathbb{P}(E|\alpha=a_{1})=1>a_{1}$. Thus, conditional on $\alpha=a_{2}$ the method is conservative and conditional on $\alpha=a_{1}$ the method is much too liberal. Thus, the method does not provide FWER control conditional on the data-dependent $\alpha$.
B.1.2 No type I error control in terms of expected discrepancy
The above raises the question question whether we at least control the FWER on average. As in the main paper, we can consider the expected value of the discrepancy ratio, which in the context of the FWER is
$\mathbb{E}(r_{\alpha}):=\mathbb{E}\big{\{}\mathbb{E}(\mathbbm{1}_{E})/\alpha|\alpha\big{\}}=\mathbb{E}(\mathbbm{1}_{E}/\alpha).$
If this value is at most 1, then we control the FWER in an average sense. However, the following proposition implies that $\mathbb{E}(\mathbbm{1}_{E}/\alpha)$ is often larger than 1. The reason that it is not always larger than 1 is due to the fact that Bonferroni is conservative under independence if $\alpha$ is fixed, i.e., when we take $a_{1}=a_{2}$.
Proposition 4.
In the setting of Proposition 3, we have $\mathbb{E}(\mathbbm{1}_{E}/\alpha)=$
$\Big{\{}1-(1-a_{2}/m)^{m}\Big{\}}a_{2}^{-1}+\Big{\{}1-(1-a_{1}/m)^{m}\Big{\}}(a_{1}^{-1}-a_{2}^{-1}),$ (8)
which can be larger than $1$, depending on $m$, $a_{1}$ and $a_{2}$. The first term in (8) is smaller than 1 and represents $\mathbb{E}(\mathbbm{1}_{E}/\alpha)$ in the situation that $a_{1}$ and $a_{2}$ are identical, i.e. the case that $\alpha$ is fixed. The second term represents the increase of $\mathbb{E}(\mathbbm{1}_{E}/\alpha)$ due to $a_{2}$ being larger than $a_{1}$.
Proof.
We have
$\mathbb{E}(\mathbbm{1}_{E}/\alpha)=\mathbb{P}\Big{(}p_{(1)}^{b}\leq a_{1}\Big{)}(1/a_{1})+\mathbb{P}\Big{(}a_{1}\leq p_{(1)}^{b}\leq a_{2}\Big{)}(1/a_{2})=$
$\mathbb{P}\Big{(}p_{(1)}\leq a_{1}/m\Big{)}(1/a_{1})+\mathbb{P}\Big{(}a_{1}/m\leq p_{(1)}\leq a_{2}/m\Big{)}(1/a_{2})=$
Like in the proof of Proposition 3, $p_{(1)}$ has a $Beta(1,m)$ distribution, so that the above equals
[…]
Recall from the proof of Proposition 3 that $B(1,m)=1/m$. Consequently, the above equals
$ma_{1}^{-1}\big{\{}-(1-x)^{m}/m\big{\}}\Big{|}_{0}^{a_{1}/m}+ma_{2}^{-1}\big{\{}-(1-x)^{m}/m\big{\}}\Big{|}_{a_{1}/m}^{a_{2}/m}=$
$ma_{1}^{-1}\Big{\{}-(1-a_{1}/m)^{m}/m--(1-0)^{m}/m\Big{\}}+ma_{2}^{-1}\Big{\{}-(1-a_{2}/m)^{m}/m--(1-a_{1}/m)^{m}/m\Big{\}}=$
$a_{1}^{-1}\Big{\{}1-(1-a_{1}/m)^{m}\Big{\}}+a_{2}^{-1}\Big{\{}(1-a_{1}/m)^{m}-(1-a_{2}/m)^{m}\Big{\}}$
$a_{1}^{-1}\Big{\{}1-(1-a_{1}/m)^{m}\Big{\}}+a_{2}^{-1}\Big{\{}(1-(1-a_{2}/m)^{m})-(1-(1-a_{1}/m)^{m})\Big{\}}=$
$\Big{\{}1-(1-a_{1}/m)^{m}\Big{\}}(a_{1}^{-1}-a_{2}^{-1})+\Big{\{}1-(1-a_{2}/m)^{m}\Big{\}}a_{2}^{-1}.$ (9)
If we plug in $m=100$, $a_{1}=0.8$ and $a_{2}=0.81$, then the above is 0.70, rounded off to two decimals. This can be understood by noting that if we take $a_{1}=a_{2}=0.8$, we actually get a similar number, since we then consider standard Bonferroni with $\alpha=0.8$ and that method is particularly conservative under independence when $\alpha$ is large.
If we plug in $a_{1}=0.05$ and $a_{2}=0.1$, then (9) is 1.44, rounded off to two decimals. This illustrates that for many values of $a_{1}$ and $a_{2}$, (9) exceeds 1. ∎
For most reasonable settings, we find with Proposition 4 that $\mathbb{E}(\mathbbm{1}_{E}/\alpha)>1$. For example, if $m=1000$, $a_{1}=0.005$ and $a_{2}=0.05$ we find $\mathbb{E}(\mathbbm{1}_{E}/\alpha)\approx 1.87.$ Thus, usually Bonferroni will too liberal, in an average sense.
B.2 Example 2: infinitely many significance levels $\alpha$
Here, we again consider $0<C\leq 1$ and assume that if $p_{(1)}^{b}\leq C$, the researcher chooses $\alpha=p_{(1)}^{b}$. If $p_{(1)}^{b}>C$, she takes $\alpha=C$. Suppose that all $m$ hypotheses are true. Then, with this definition of $\alpha$, there is always at least one incorrect rejection if $p_{(1)}^{b}\leq C$. Thus, conditional on $\alpha<C$, $\mathbbm{1}_{E}$ is always equal to 1. Hence, we clearly do not have $\mathbb{E}(\mathbbm{1}_{E}/\alpha|\alpha<C)\leq 1$, i.e., we do not have FWER control conditional on $\alpha<C$.
Next we can investigate whether we at least control the FWER on average in the sense that $\mathbb{E}(r_{\alpha})=\mathbb{E}(\mathbbm{1}_{E}/\alpha)\leq 1.$ The following proposition states that this is not the case. In fact, $\mathbb{E}(\mathbbm{1}_{E}/\alpha)$ is infinity, like we saw in the case of one hypothesis.
Proposition 5.
Consider the setting of Propositions 3 and 4 above, but define $\alpha=p_{(1)}^{b}$ if $p_{(1)}^{b}\leq C$ and $\alpha=C$ otherwise. Then
$\mathbb{E}(r_{\alpha})=\mathbb{E}(\mathbbm{1}_{E}/\alpha)=\infty.$
Proof.
We will again use that $p_{(1)}$ has a $Beta(1,m)$ distribution, so has density function $f(x)=B(1,m)(1-x)^{m-1}$. Note that $E$ happens if and only if $p_{(1)}^{b}\leq C$, i.e., if and only if $p_{(1)}\leq C/m$. Hence, $\mathbb{E}(\mathbbm{1}_{E}/\alpha)=$
[…]
where we used that if $p_{(1)}^{b}\leq C$ then $1/\alpha=1/(mp_{(1)})$. Since $B(1,m)^{-1}=m$, the above equals
[…]
Note that $(1-x)^{m-1}\geq(1-C/m)^{m-1}$ for all $0\leq x\leq C/m$. Hence the above is at least
[…]

### v2

@@
 1.1 The call to decrease significance thresholds
-Recently, there has been an intense debate in many scientific fields on whether the significance threshold for declaring scientific discoveries should be changed from $0.05$ to $0.005$ (Borenstein, 2001; Ioannidis, 2005; Banerjee et al., 2009; Mudge et al., 2012; Martin and Bridgmon, 2012; Kim and Ji, 2015; Ioannidis, 2016; Szucs and Ioannidis, 2017; Benjamin et al., 2018; Ioannidis, 2018; Held, 2019; Di Leo and Sardanelli, 2020; Khan et al., 2020; Aguinis et al., 2021; Schnog et al., 2021; Thakur and Jha, 2022; Maier and Lakens, 2022; Fitzpatrick et al., 2024).
-In most statistics textbooks, the “significance threshold” or “significance level” is denoted by $\alpha$ and is assumed to be prespecified, i.e., data-independent (Van der Vaart, 1998; Lehmann and Romano, 2022; Hansen, 2022).
-In particular, $\alpha$ is assumed independent of the p-value, since otherwise $\alpha$ would depend on the data (Hubbard and Bayarri, 2003).
-One important paper that stimulated the mentioned debate, bears the title “Redefine statistical significance” (Benjamin et al., 2018).
-In this article, 72 influential scientists propose changing the “threshold for statistical significance for claims of new discoveries” to $0.005$.
-As a consequence of this proposal, scientists may find themselves in a field where there is not one standard threshold, but multiple: some groups of scientists may support $\alpha=0.005$ and some $\alpha=0.05$.
+Conducting a hypothesis test requires choosing a “significance level” or “significance threshold”, typically denoted by $\alpha$ [48, 32, 13].
+It is assumed to be prespecified, and therefore independent of the p-value [17].
+Recently, there has been an intense debate in many fields on whether the threshold should be changed from 0.05 to 0.005 [7, 20, 3, 37, 34, 26, 21, 43, 4, 22, 29, 14, 8, 25, 1, 41, 44, 9].
+In one important paper that stimulated the debate, 72 influential scientists propose changing the “threshold for statistical significance for claims of new discoveries” to 0.005 [4].
+As a consequence of such proposals, scientists may find themselves in a field where there is not one standard threshold, but multiple: some groups of scientists may support $\alpha=0.005$ and some $\alpha=0.05$.
 This can also translate into the implicit or explicit preferences of journals, with some journals preferring $0.005$ and others $0.05$.
 1.2 The statistical danger of differing thresholds
-The message of this paper is that if within one field, there are multiple standard significance thresholds, then researchers may be tempted to first analyze their data and then choose the $\alpha$ that suits them.
-The problem can best be indicated with an example.
-Suppose a researcher is interested in a certain null hypothesis and plans to compute a p-value for it.
-In her field there are two standard significance thresholds, $0.005$ and $0.05$, simply because some researchers or journal editors in her field advocate 0.005 and others advocate 0.05.
+The message of this paper is that if within one field, there are multiple standard significance thresholds, then researchers may be tempted to first analyze their data and then choose the threshold $\alpha$ that suits them.
+For example, suppose a researcher is interested in a certain null hypothesis and plans to compute a p-value for it.
+In her field there are two standard significance thresholds, $0.005$ and $0.05$, because some researchers or journal editors in her field advocate 0.005 and others advocate 0.05.
 After she has computed the p-value, she wants to share it in some form, e.g.
@@
 Then the choice of where to present or where to submit her results will potentially depend on the p-value that she finds.
-For example, if she finds $p=0.02$, then she may be more likely to submit to a journal that accepts $\alpha=0.05$ than to a journal that prefers $\alpha=0.005$.
-On the other hand, if she finds an extremely small p-value, then she may prefer a journal with $\alpha=0.005$, since then her finding is significant at a more stringent threshold.
+For example, if she finds $p=0.003$, then she may prefer a journal that supports $\alpha=0.005$ over a journal that supports $\alpha=0.05$, since then her finding is significant at a more stringent threshold.
 Consequently, it would be tempting for the researcher not to fix her $\alpha$ before running her experiment, but to let her $\alpha$ depend on the journal or conference, which she chooses after obtaining her p-value.
 The statistical issue that arises, is that the researcher’s $\alpha$ is no longer independent of her p-value.
-Consequently, the test loses its usual decision-theoretic interpretation, i.e., there is no guarantee anymore that the type I error probability is below $\alpha$ (Hubbard and Bayarri, 2003; Grünwald, 2024; Koning, 2024a).
+Consequently, the test loses its usual decision-theoretic interpretation, i.e., there is no guarantee anymore that the type I error probability is below $\alpha$ [17, 12, 28].
 Note that this issue is about more than publication bias.
 It is about the validity of an individual test, from the perspective of the individual researcher: when she lets $\alpha$ depend on the data, she cannot validly interpret her own statistical test anymore.
-In theory, a solution would be that the researcher fixes her significance threshold beforehand and never changes it.
-However, one can easily imagine a researcher who finds $p=0.003$ and then claims her significance level is $\alpha=0.005$, even though she never fixed $\alpha$ beforehand.
+In theory, a solution would be that the researcher fixes her significance threshold beforehand, but one can easily imagine a researcher who finds $p=0.003$ and then claims her significance level is $\alpha=0.005$, even though she never fixed $\alpha$ beforehand.
 1.3 A related discussion: abandoning significance thresholds
-In parallel to the discussion on what the threshold should be, there has been a discussion on whether significance thresholds should be used at all (Hubbard and Lindsay, 2008; Wasserstein and Lazar, 2016; McShane et al., 2019; Amrhein et al., 2019).
-Indeed, a downside of significance testing in practice is that it may overemphasize a specific p-value threshold rather than interpreting p-values continuously, considering effect sizes, and using other scientifically relevant information (Borenstein, 1994; Greenland et al., 2016; Thompson, 2007; Betensky, 2019; Scheel et al., 2021; Imbens, 2021; Muff et al., 2022).
-For example, the famous ASA statement on p-values (Wasserstein and Lazar, 2016) notes that “Scientific conclusions and business or policy decisions should not be based only on whether a p-value passes a specific threshold.” Note that the ASA statement is not against p-values, but against too much focus on p-value thresholds.
-Many statisticians agree with the ASA statement’s point, but nevertheless we see that in the recent statistical literature, significance testing is still prominent (Lehmann and Romano, 2022; Hansen, 2022).
+In parallel to the discussion on what the threshold should be, there has been a discussion on whether significance thresholds should be used at all [18, 50, 36, 2].
+Indeed, a downside of significance testing in practice is that it may overemphasize a specific p-value threshold rather than interpreting p-values continuously, considering effect sizes, and using other scientifically relevant information [6, 10, 45, 5, 40, 19, 38].
+For example, the famous ASA statement on p-values [50] notes that “Scientific conclusions and business or policy decisions should not be based only on whether a p-value passes a specific threshold.” Note that the ASA statement is not against p-values, but against too much focus on p-value thresholds.
+Many statisticians agree with the ASA statement’s point, but nevertheless we see that in the recent statistical literature, significance testing is still prominent [32, 13].
 One reason is that from a mathematical point of view, there is often essentially no difference between a formula for a significance test, a formula for a p-value and a formula for a confidence interval.
@@
 Another reason is that some statisticians still value type I error control, which is only possible if some $\alpha$ is chosen.
-Hence, although many statisticians largely agree with the ASA statement, they continue to develop hypothesis tests.
-Beyond theoretical statistics, in science in general we see that significance testing has not disappeared (Hubbard, 2019; Mayo and Hand, 2022; Thakur and Jha, 2022; Maier and Lakens, 2022; Hansen, 2022; Fitzpatrick et al., 2024).
+Beyond theoretical statistics, in science in general we see that significance testing has not disappeared [16, 35, 44, 33, 13, 9].
 Moreover, if one abandons significance tests, then one is still likely to compute confidence intervals, which also require choosing $\alpha$.
-In that case, choosing $\alpha$ post hoc causes issues similar to those discussed in this paper.
 1.4 This paper
 Throughout the paper, we stay within the classical Neyman-Pearson theory of hypothesis testing.
-We note that one of the main reasons why classical Neyman-Pearson hypothesis testing works in practice, is that there exists a single standard level $\alpha$ in most fields.
-This automatically avoids that $\alpha$ is chosen based on the data.
-Although it is known that having $\alpha$ depend on the p-value is invalid (Hubbard and Bayarri, 2003), this paper seems to be the first to point out this will likely occur when within a field there are multiple standard thresholds.
+Our main point is that tests are often invalidated if $\alpha$ depends on the p-value [17], and that this will likely occur when within a field there are multiple standard thresholds.
+[46] make the following related point: one possible reason why a single $\alpha$ has been the norm for so long, is that it prevents researchers from choosing $\alpha$ post hoc.
 Further, we discuss in detail what can go wrong in terms of type I error control.
 We evaluate what the type I error rate is, conditional on a certain $\alpha$ being chosen based on the p-value.
+Moreover, we introduce a measure of how problematic a test with data-dependent $\alpha$ is.
+If this measure is large, this means the conditional type I error rate of the test is too large in an average sense.
 Apart from discussing the situation where the researcher chooses between two significance levels, we will also consider the most extreme scenario where there are many choices for $\alpha$.
-Moreover, for a test with a data-dependent level $\alpha$, we consider a measure of evidence for how problematic this is on average.
-If this measure is large, this means the conditional type I error rate of the test is much too large in an average sense.
-In the examples that we provide in Sections 3 and 4, the interpretation of the hypothesis test is ruined due to the post hoc selection of $\alpha$.
-Throughout, we focus on the case where one hypothesis is tested.
-In case there are multiple hypotheses and a multiple testing correction is used, similar issues arise when $\alpha$ is chosen post hoc.
-This is discussed in the Supplementary Material.
 In Section 3 and the Discussion, we briefly point out some potential solutions to the discussed issue.
-2 Notation and concepts
-2.1 Tests with data-dependent $\alpha$
-Consider any hypothesis $H_{0}$ and a corresponding hypothesis test $\phi_{\alpha}$.
-The subscript is used because the test depends on the chosen significance level $\alpha>0$.
-The test $\phi_{\alpha}$ is a function from the sample space to the set $\{0,1\}$, where $1$ indicates that $H_{0}$ is rejected and 00 indicates that $H_{0}$ is not rejected.
-We will indicate the data with $X$, which takes values in the sample space.
-We view the test $\phi_{\alpha}=\phi_{\alpha}(X)$ as a function of $X$, i.e., given $\alpha$, the outcome of the test $\phi_{\alpha}$ is entirely determined by the data $X$.
-In classical hypothesis testing, the level $\alpha$ is always assumed to be selected independently from the data.
-For such an independently selected $\alpha$, it is typical to use a test $\phi_{\alpha}$ with the guarantee that it does not reject the hypothesis too often when it is true, namely with probability at most $\alpha$.
-We assume throughout that if $\alpha$ is chosen independently from $X$, then the tests we consider are exactly valid, in the sense that $\mathbb{P}(\phi_{\alpha}=1)=\alpha$ if $H_{0}$ is true.
-As indicated in the introduction, there are scenarios where $\alpha$ would depend on $X$ – for example if $\alpha$ depends on the p-value.
-Then $\alpha$ is a function of $X$, although we will often write $\alpha$ instead on $\alpha(X)$ for short.
-If $\alpha$ is data-dependent, then conditional on $\alpha$ taking a certain value, we possibly lose the guarantee that $\mathbb{P}_{H_{0}}(\phi_{\alpha}=1)\leq\alpha$.
-For example, suppose $\alpha$ takes two possible values, e.g.
-0.005 and 0.05, depending on the data $X$.
-Then, depending on how $\alpha$ exactly depends on $X$, it can be the case that $\mathbb{P}({\phi_{\alpha}=1|\alpha=0.005})$ is not at most $0.005$, but larger than $0.005$.
-This will be illustrated in Section 3.
-2.2 Measuring the discrepancy between $\alpha$ and the type I error rate
-Let $A\subseteq[0,1]$ denote the set of possible values that $\alpha(X)$ can take.
+2 Discrepancies between $\alpha$ and the type I error rate
+2.1 Notation and concepts
+Consider a null hypothesis $H_{0}$ and a corresponding test $\phi_{\alpha}$, where $0<\alpha\leq 1$ indicates the chosen significance level.
+The test $\phi_{\alpha}=\phi_{\alpha}(X)$ is a function of the data $X$ and maps to the set $\{0,1\}$, where $1$ indicates that $H_{0}$ is rejected.
+In this paper, all mentioned probabilities and expectations will be under $H_{0}$.
+Suppose that $\alpha$ possibly depends on $X$.
+Let $A\subseteq[0,1]$ denote the set of possible values that $\alpha$ can take.
 Throughout we use ‘$a$’ to denote a value from $A$, independent of the data.
 We can then ask whether conditional on the data-dependent $\alpha$ taking the value $a$, the type I error rate is below $\alpha$, i.e.
-whether $\mathbb{P}_{H_{0}}(\phi_{\alpha}=1|\alpha=a)\leq a$.
-A measure of the discrepancy between $a$ and $\mathbb{P}_{H_{0}}(\phi_{\alpha}=1|\alpha=a)$ is the difference
-$d_{a}:=\mathbb{P}\big{(}\phi_{\alpha}=1|\alpha=a\big{)}-a.$ (1)
-Ideally, we would have $d_{a}=0$ for all $a\in A$, but this is often not the case for many $a\in A$, if $\alpha$ depends on the data.
-This is well known (Hubbard and Bayarri, 2003) and illustrated in Sections 3 and 4.
+whether $\mathbb{P}(\phi_{\alpha}=1\mid\alpha=a)\leq a$.
+A measure of the discrepancy between $a$ and $\mathbb{P}(\phi_{\alpha}=1\mid\alpha=a)$ is the difference
+$d_{a}:=\mathbb{P}\big{(}\phi_{\alpha}=1\mid\alpha=a\big{)}-a.$
+Ideally, we would have $d_{a}=0$ for all $a\in A$, but this is often not the case for many $a\in A$, if $\alpha$ depends on the data. This is well known [17] and illustrated in Sections 2.2 and 2.3.
 Now suppose that for some $a\in A$, the difference is $d_{a}=0.01$.
@@
 To a large extent, this depends on what $a$ is.
-For example, if $a=0.05$, then having $d_{a}=0.01$ is not highly problematic, since $\mathbb{P}(\phi_{\alpha}=1|\alpha=a)=0.06$ is then not much larger than $a$, in relative terms.
-However, if $a=0.005$, then $d_{a}=0.01$ implies that the true error rate $\mathbb{P}_{H_{0}}(\phi_{\alpha}=1|\alpha=a)=0.015$ is three times larger than $a$.
+For example, if $a=0.05$, then having $d_{a}=0.01$ is not highly problematic, since $\mathbb{P}(\phi_{\alpha}=1\mid\alpha=a)=0.06$ is then not much larger than $a$, in relative terms.
+However, if $a=0.005$, then $d_{a}=0.01$ implies that the true error rate $\mathbb{P}(\phi_{\alpha}=1\mid\alpha=a)=0.015$ is three times larger than $a$.
 In that case, $d_{a}=0.01$ is a huge difference.
@@
 Instead it can be more meaningful to look at the relative discrepancy
-$r_{a}:=\mathbb{P}_{H_{0}}\big{(}\phi_{\alpha}=1|\alpha=a\big{)}/a,$ (2)
-where we remind the reader that $\alpha=\alpha(X)$ depends on the data.
+$r_{a}:=\mathbb{P}\big{(}\phi_{\alpha}=1\mid\alpha=a\big{)}/a.$
 We will call $r_{a}$ the discrepancy ratio conditional on $\alpha=a$.
 We would like this ratio to be at most 1.
-If $r_{a}<1$, then the test is conservative conditional on $\alpha=a$, and if $r_{a}>1$ the test is liberal conditional on $\alpha=a$.
+If $r_{a}>1$, the test is liberal conditional on $\alpha=a$.
 Note that for a given number $a$, $r_{a}$ is a fixed number.
-In contrast, since $\alpha=\alpha(X)$ depends on the data, $r_{\alpha}$ is data-dependent.
-2.3 Expected discrepancy
-If the discrepancy ratio $r_{a}$ is below 1 for every $a\in A$, then in particular the expected value of the discrepancy ratio will be below 1, i.e.,
-$\mathbb{E}r_{\alpha}=\mathbb{E}\big{\{}\mathbb{P}(\phi_{\alpha}=1|\alpha)/\alpha\big{\}}$ (3)
-will be at most 1.
-Note that this is the expected value of a conditional probability divided by $\alpha$.
-Likewise, if (3) is larger than 1, it follows that there is at least one $a\in A$ for which (2) is larger than 1.
-Note that in general, if $\mathbb{E}r_{\alpha}$ is much larger than $1$ under $H_{0}$, it follows that either there is at least a small probability that $r_{\alpha}$ is much larger than $1$, or there is a large probability that $r_{\alpha}$ is at least somewhat larger than 1.
+In contrast, since $\alpha$ depends on the data, $r_{\alpha}$ is data-dependent.
+If the discrepancy ratio $r_{a}$ is below 1 for every $a\in A$, then in particular $\mathbb{E}r_{\alpha}$ will be below 1.
+If on the other hand $\mathbb{E}r_{\alpha}$ is much larger than $1$ under $H_{0}$, it follows that either there is at least a small probability that $r_{\alpha}$ is much larger than $1$, or there is a large probability that $r_{\alpha}$ is at least somewhat larger than 1.
 Both these possibilities are problematic for researchers.
@@
 If $\mathbb{E}r_{\alpha}\gg 1$, this indicates that the test procedure is seriously flawed, in an average sense.
-It is useful to note that $\mathbb{E}r_{\alpha}$ can be rewritten in the following ways:
-$\mathbb{E}r_{\alpha}=\mathbb{E}\big{\{}\mathbb{P}(\phi_{\alpha}=1|\alpha)/\alpha\big{\}}=\mathbb{E}\big{\{}\mathbb{E}\big{(}\phi_{\alpha}|\alpha\big{)}/\alpha\big{\}}=$
-$\mathbb{E}\big{\{}\mathbb{E}\big{(}\phi_{\alpha}/\alpha|\alpha\big{)}\big{\}}=\mathbb{E}\big{(}\phi_{\alpha}/\alpha\big{)}.$ (4)
-The equalities (4) tell us that $\mathbb{E}r_{\alpha}$ can be rewritten as the expectation of a conditional expectation, which can in turn be rewritten as a simple expectation.
-3 Example 1: two possible significance levels $\alpha$
-Consider a researcher who tests one hypothesis.
-In her field there are two standard significance thresholds, $a_{1}$ and $a_{2}$, where $10<a_{1}<a_{2}<$ 1.
-As explained in the Introduction, there are scenarios where her $\alpha$ would depend on the p-value.
-In particular, if the researcher obtains a p-value $p$ with $a_{1}<p\leq a_{2}$, she may be more likely to take $\alpha=a_{2}$ than when $p\leq a_{1}$, in which case she could be more likely to take $\alpha=a_{2}$.
-As a simple behavioural model, suppose that the researcher takes $\alpha=a_{1}$ if $p\leq a_{1}$ and takes $\alpha=a_{2}$ if $a_{1}<p\leq 1$.
-We will use the notation from Section 2.1.
-Since the threshold depends on the p-value, it depends on the data.
-Thus, $\alpha(X)$ depends on $X$, although we supress ‘$(X)$’ for brevity.
-We have $\phi_{\alpha}=\phi_{\alpha(X)}=1$ if and only if $p\leq\alpha$.
-If $\alpha$ would be fixed, then the test $\phi_{\alpha}$ would be valid in the sense that $\mathbb{P}(\phi_{\alpha}=1)\leq\alpha$.
-However, because $\alpha$ depends on the p-value, the interpretation of the test is ruined, as we will see.
-3.1 No type I error control conditional on $\alpha$ or marginally
-We investigate the type I error probability conditional on a certain value for $\alpha$ being chosen.
-What happens conditional on $\alpha=a_{1}$ can be seen immediately: if $\alpha=a_{1}$, then this means that $p$ was apparently at most $a_{1}$.
-This means that $H_{0}$ is rejected.
-Thus, under $H_{0}$, conditional on $\alpha=a_{1}$, the type I error probability is not $a_{1}$, but 1.
-Further, the type I error probability is larger than $a_{1}$ marginally.
-Thus, the familiar interpretation of the hypotheses test is ruined.
-This means that the researcher cannot validly interpret her own statistical test anymore, in the usual decision-theoretic way.
-The following proposition states what happens to the type I error probability conditional on $\alpha=a_{2}$. The proof is in the appendix.
-Proposition 1.
-For the above procedure that takes $\alpha=a_{1}$ if $p\leq a_{1}$ and $\alpha=a_{2}$ otherwise, under $H_{0}$ we have
-$\mathbb{P}(\phi_{\alpha}=1|\alpha=a_{2})=\big{(}(1-a_{1}/a_{2})/(1-a_{1})\big{)}a_{2}<a_{2}.$
-Thus, conditional on the data-dependent $\alpha$ being $a_{2}$, the type I error rate does not exceed $a_{2}$, although the test is conservative since $\mathbb{P}(\phi_{\alpha}=1)$ is strictly smaller than $a_{2}$.
-For example, for $a_{1}=0.005$ and $a_{2}=0.05$, we have $\mathbb{P}(\phi_{\alpha}=1|\alpha=a_{2})=a_{2}((1-0.005/0.05)/(1-0.005))\approx a_{2}\cdot 0.9045$.
-Thus, in the notation of Section 2.2, $r_{a_{2}}\approx 0.9045$.
-In summary, conditional on $\alpha=a_{1}$ the type I error probability is too large and conditional on $\alpha=a_{2}$ the type I error probability is on the conservative side.
-Put differently, the discrepancy ratio is $r_{\alpha}>1$ if $\alpha=a_{1}$ and $r_{\alpha}<1$ if $\alpha=a_{2}$.
-3.2 No type I error control in terms of expected discrepancy
-Next, we can ask whether perhaps we control the type I error probability in expectation, i.e., whether $\mathbb{E}r_{\alpha}\leq 1$ under $H_{0}$.
-This would mean that at least in an average sense, we have type I error control.
-The following proposition implies that this does not hold.
-Proposition 2.
-For the procedure that takes $\alpha=a_{1}$ if $p\leq a_{1}$ and $a_{2}$ otherwise, under $H_{0}$ we have
-$\mathbb{E}r_{\alpha}=1+(a_{2}-a_{1})/a_{2}>1.$
-For example, if $a_{1}=0.005$ and $a_{2}=0.05$, then $\mathbb{E}_{H_{0}}(r_{\alpha})=1.9$.
-Since $\mathbb{E}_{H_{0}}(r_{\alpha})$ is not bounded by 1, it follows that $r_{\alpha}$ is larger than 1 conditional on at least one possible value of $\alpha$.
-We already saw that this is the case for $\alpha=a_{1}$.
-We conclude that both conditionally and on average, the researcher’s test does not provide type I error control.
-4 Example 2: Infinitely many significance levels $\alpha$
-Now suppose there are not two thresholds to choose from but many.
-To take this to the extreme, suppose each $0<\alpha\leq C$ can be chosen.
-Here $0<C\leq 1$ may for example be $0.1$.
-It is then easily imaginable that the author is biased towards choosing an $\alpha$ that is at least as large as her p-value $p$, when possible.
-To fix ideas, suppose the researcher in fact selects $\alpha=p$, unless $p>C$, in which case she selects $\alpha=C$.
-As usual, we denote the test by $\phi_{\alpha}$, where $\alpha$ depends on $p$.
-In case $p\leq C$, we always have $p\leq\alpha$, so that $H_{0}$ is always rejected.
-Likewise, if $\alpha<C$, $H_{0}$ is always rejected.
-Thus, under $H_{0}$, conditional on $\alpha<C$, the type I error rate is not below $\alpha$.
-Specifically, for every $a<C$, we have $r_{a}=\mathbb{E}(\phi_{\alpha}|\alpha=a)/a=1/a$.
-Thus, the smaller $a$ is, the larger $r_{a}$ is, i.e.
-the larger discrepancy ratio is conditional on $\alpha=a$.
-Next, we may ask whether we have type I error control not conditionally, but at least on average, in the sense that $\mathbb{E}(r_{\alpha})\leq 1$ under $H_{0}$.
+It is useful to note that $\mathbb{E}r_{\alpha}$ can be rewritten as a simple expectation:
+$\mathbb{E}r_{\alpha}$ $=\mathbb{E}\big{\{}\mathbb{P}(\phi_{\alpha}=1\mid\alpha)/\alpha\big{\}}=\mathbb{E}\big{\{}\mathbb{E}\big{(}\phi_{\alpha}|\alpha\big{)}/\alpha\big{\}}$
+$=\mathbb{E}\big{\{}\mathbb{E}\big{(}\phi_{\alpha}/\alpha\mid\alpha\big{)}\big{\}}=\mathbb{E}\big{(}\phi_{\alpha}/\alpha\big{)}.$ (1)
+2.2 Example 1: two possible significance levels $\alpha$
+Consider a researcher who tests one hypothesis. She computes a p-value, which we assume to exactly satisfy
+$\mathbb{P}(p\leq\alpha)=\alpha$
+under $H_{0}$ for every data-independent $\alpha$.
+Suppose that in her field there are two standard significance thresholds, $a_{1}$ and $a_{2}$, where $10<a_{1}<a_{2}<$ 1.
+As a simple behavioral model, assume she chooses the smaller threshold $\alpha=a_{1}$ if $p\leq a_{1}$ and chooses $\alpha=a_{2}$ if $a_{1}<p\leq 1$, so that $\alpha$ depends on $p$ and hence on the data.
+What happens to the type I error probability conditional on $\alpha=a_{1}$ can be seen immediately: if $\alpha=a_{1}$, then this means that $p$ was apparently at most $a_{1}$, which means that $H_{0}$ is rejected.
+Thus, under $H_{0}$, conditional on $\alpha=a_{1}$, the type I error probability is not $a_{1}$ but 1.
+Conditional on $\alpha=a_{2}$, the type I error probability is strictly smaller than $a_{2}$,
+$\mathbb{P}(p\leq\alpha\mid\alpha=a_{2})$ $=\mathbb{P}(p\leq a_{2}\mid p>a_{1})$
+$=\frac{a_{2}-a_{1}}{1-a_{1}}<\frac{a_{2}-a_{2}a_{1}}{1-a_{1}}=a_{2}.$
+Thus, the discrepancy ratio is $r_{\alpha}>1$ if $\alpha=a_{1}$ and $r_{\alpha}<1$ if $\alpha=a_{2}$.
+Next, we can ask whether perhaps we control the type I error probability in expectation, in the sense that $\mathbb{E}r_{\alpha}\leq 1$ under $H_{0}$.
+This turns out to not be the case either, since under $H_{0}$, by (1),
+$\mathbb{E}r_{\alpha}$ $=\mathbb{E}(\phi_{\alpha}/\alpha)$
+$=\mathbb{P}(p\leq a_{1})\mathbb{E}(\phi_{\alpha}/\alpha\mid p\leq a_{1})$
+$+\mathbb{P}(a_{1}<p\leq a_{2})\mathbb{E}(\phi_{\alpha}/\alpha\mid a_{1}<p\leq a_{2})$
+$=a_{1}(1/a_{1})+(a_{2}-a_{1})(1/a_{2})$
+$=1+(a_{2}-a_{1})/a_{2}>1.$
+For example, if $a_{1}=0.005$ and $a_{2}=0.05$, then $\mathbb{E}r_{\alpha}=1.9$. We conclude that both conditionally and on average, the researcher’s test does not provide type I error control.
+2.3 Example 2: Infinitely many significance levels $\alpha$
+Now suppose there are not just two, but many thresholds to choose from.
+In fact, let us take this to the extreme by allowing the researcher to choose any threshold $0<\alpha\leq C$ up to some constant $C>0$, e.g.
+$C=0.05$ or $C=0.1$.
+If the researcher desires to report a discovery, it is easy to imagine that she is biased towards choosing an $\alpha$ that is at least as large as her p-value $p$, when possible.
+Again taking the most extreme choice, suppose she selects $\alpha=p$, unless $p>C$, in which case she selects $\alpha=C$.
+In case $\alpha<C$, $H_{0}$ is always rejected, so that we do not have conditional type I error rate control.
+Indeed, for every $a<C$, we have $r_{a}=1/a>1$.
+Next, we may ask whether we have type I error control not conditionally, but at least on average, in the sense that $\mathbb{E}r_{\alpha}\leq 1$ under $H_{0}$.
 Since $p$ is uniform on $[0,1]$ and $p\leq\alpha$ if and only if $p\leq C$, we have
-$\mathbb{E}(r_{\alpha})=\mathbb{E}(\phi_{\alpha}/\alpha)=$
-$\mathbb{P}(p>\alpha)\cdot 0+\mathbb{P}(p\leq\alpha)\mathbb{E}(\phi_{\alpha}/\alpha|p\leq\alpha)=$
-$\mathbb{P}(p\leq\alpha)\mathbb{E}(\phi_{\alpha}/\alpha|\alpha=p\leq C)=$
-$\mathbb{P}(p\leq\alpha)\mathbb{E}(1/p|p\leq C)=$
+$\mathbb{E}r_{\alpha}$ $=\mathbb{E}(\phi_{\alpha}/\alpha)$
+$=\mathbb{P}(p>\alpha)\cdot 0+\mathbb{P}(p\leq\alpha)\mathbb{E}(\phi_{\alpha}/\alpha\mid p\leq\alpha)$
+$=\mathbb{P}(p\leq\alpha)\mathbb{E}(\phi_{\alpha}/\alpha\mid\alpha=p\leq C)$
+$=\mathbb{P}(p\leq\alpha)\mathbb{E}(1/p\mid p\leq C)$
 […]
-We see that the expected discrepancy ratio $\mathbb{E}(r_{\alpha})$ is not only larger than 1, but in fact infinity. Thus, in expectation, the test is completely out of control.
+We see that the expected discrepancy ratio $\mathbb{E}r_{\alpha}$ is not only larger than 1, but in fact infinity. Thus, in expectation, the test is completely out of control.
 3 Connection to e-values
-A partial solution to the problem of data-dependent $\alpha$ are so-called e-values.
-An e-value is a measure of evidence that has been recently proposed as an alternative to the p-value (Howard et al., 2021; Shafer, 2021; Vovk and Wang, 2021; Grünwald et al., 2024; Ramdas et al., 2023; Koning, 2024b).
-An e-value is typically defined as a non-negative random variable $e$, whose expectation is bounded by 1 under the hypothesis $H_{0}$:
-$\mathbb{E}\,e\leq 1.$ (5)
-In the context of a simple null and alternative hypothesis that only contain a single distribution, a prominent example of an $e$-value is the likelihood ratio between these distributions (Grünwald et al., 2024; Larsson et al., 2024).
+In case we are only interested in $\mathbb{E}r_{\alpha}$, then a solution is offered by the e-value.
+An e-value is a measure of evidence that has been recently proposed as an alternative to the p-value [15, 42, 49, 11, 39, 27].
+An e-value is typically defined as a non-negative random variable $e$, whose expectation is bounded by 1 under $H_{0}$:
+$\mathbb{E}\,e\leq 1.$
+In the context of a simple null and alternative hypothesis that only contain a single distribution, a prominent example of an $e$-value is the likelihood ratio between these distributions [42].
+For a general null, desirable e-values are increasing functions of the likelihood ratio between the alternative and a kind of least-favorable distribution in the null [11, 31, 27].
 The reciprocal $p^{*}=1/e$ of an e-value gives rise to a conservative p-value. Indeed, every such $p$-value is valid:
 $\mathbb{P}(p^{*}\leq\alpha)=\mathbb{P}(e\geq 1/\alpha)\leq\alpha\,\mathbb{E}\,e\leq\alpha,$
-for all $\alpha>0$, where the first inequality follows from Markov’s inequality and the second inequality from the definition of the e-value.
-Recently, Koning (2024a) has shown that such a p-value is not just some conservative p-value, but has deep links to the data-dependent selection of the level and the expected relative discrepancy (see also Grünwald, 2024).
-In particular, an e-value-based p-value $p^{*}$ leads to decisions that have an expected discrepancy ratio below 1 for every data-dependent choice of the level $\alpha$.
-That is, for every possibly data-dependent level $\alpha$,
-$\mathbb{E}\,r_{\alpha}=\mathbb{E}\left[\frac{\mathbb{P}(p^{*}\leq\alpha)}{\alpha}\right]\leq 1.$ (6)
-A consequence is that a test $\phi_{\alpha}=\mathbb{I}\{p^{*}\leq\alpha\}$ based such a special p-value $p^{*}$ is valid for any data-dependent $\alpha$, in the sense that the expected relative discrepancy remains below 1.
-This is in stark contrast to traditional p-values, for which this error can be $\infty$, as discussed in Section 4.
-The connection is that a traditional $p$-value only satisfies (6) for every data-independent level.
-While e-values are useful when $\alpha$ depends on the data, they should not be used when $\alpha$ is already fixed in advance.
-Indeed, in that case there is a considerable power loss, compared to using a traditional p-value.
+for all data-independent $\alpha>0$, where the first inequality follows from Markov’s inequality and the second inequality from the definition of the e-value.
+Recently, [28] showed that $p^{*}$ is not just some overly conservative p-value, but exactly the kind of p-value that satisfies
+$\mathbb{E}\,r_{\alpha}=\mathbb{E}\left(\frac{\mathbb{P}(p^{*}\leq\alpha\mid\alpha)}{\alpha}\right)\leq 1$
+regardless of how $\alpha$ depends on the data. This is in stark contrast to traditional p-values, for which $\mathbb{E}r_{\alpha}$ can be $\infty$, as discussed in Section 2.3.
 Discussion
 The fundamental point of this paper is that significance testing works in practice because researchers all by default use the same $\alpha$.
-This ensures that $\alpha$ is independent of the data.
+This ensures that $\alpha$ is independent of the data [46].
 If there are multiple standards for $\alpha$ in a field and a researcher chooses her $\alpha$ post hoc, then that spells the end of the validity of her tests.
-Her p-values may be perfectly valid and they will provide a pragmatic measure of the amount of proof against $H_{0}$ (Efron, 2010, p.31), but she will not have a test with the usual interpretation in terms of type I error control.
 Likewise, confidence intervals are no longer valid if $\alpha$ is chosen post hoc.
-As statisticians we should give this some careful thought.
-One solution would be that researchers fix $\alpha$ in advance and never change it, but they would not necessarily have enough incentives or awareness to do this.
+One solution would be that researchers fix $\alpha$ in advance and do not change it, but they would not necessarily have enough incentives or awareness to do this.
 When a community of researchers within a scientific field starts adopting a lower $\alpha$, say $\alpha=0.005$, there are two possible scenarios: 1.
@@
 The problem described in this paper occurs in scenario 2.
-Existing papers that advocate decreasing $\alpha$ (Johnson et al., 2017; Szucs and Ioannidis, 2017; Benjamin et al., 2018; Ioannidis, 2018; Held, 2019; Schnog et al., 2021; Fitzpatrick et al., 2024), do not discuss the following question: would they already consider it an improvement if some people or journals in a field decrease $\alpha$, or do they consider it essential that the field as a whole uses the same $\alpha$?
+Existing papers that advocate decreasing $\alpha$ [23, 43, 4, 22, 14, 41, 9], do not discuss the following question: would they already consider it an improvement if some people or journals in a field decrease $\alpha$, or do they consider it essential that the field as a whole uses the same $\alpha$?
 The present paper illustrates that these distinctions matter.
-Suppose that in a field the scenario occurs where some stick to $\alpha=0.05$ while others adopt $\alpha=0.005$ – thus providing an incentive for letting $\alpha$ depend on the data.
-One solution would be for journals to require pre-registration of the $\alpha$ before researchers conduct an experiment.
+Suppose that in a field the scenario occurs where some stick to $\alpha=0.05$ while others adopt $\alpha=0.005$—thus providing an incentive for letting $\alpha$ depend on the data.
+One solution would be to require pre-registration of $\alpha$ before researchers conduct an experiment [33].
 Another partial solution would be to use e-values instead of p-values, as discussed in Section 3.
-Another way out is that a researcher uses a Bayesian approach, which does not require fixing an $\alpha$ and nevertheless tells us the probabilty that $H_{0}$ is false – which frequentist statistics never does (Kelter, 2020; Lakens et al., 2020; van de Schoot et al., 2021).
+Another way out is that a researcher uses a Bayesian approach, which does not require fixing an $\alpha$ and nevertheless tells us the probabilty that $H_{0}$ is false – which frequentist statistics never does [24, 30, 47].
 A potential downside of Bayesian inference is that one needs to specify a prior distribution, which often depends on subjective considerations.
-Appendix A Proofs
-A.1 Proof of Proposition 1
-Proof.
-Since $\alpha=a_{2}\Leftrightarrow p>a_{1}$ and $p\leq a_{2}\Leftrightarrow\phi_{\alpha}=1$, under $H_{0}$ we have
-$\mathbb{P}(\phi_{\alpha}=1|\alpha=a_{2})=$
-$\mathbb{P}\big{\{}p\leq a_{2}|p>a_{1}\big{\}}=$
-$\mathbb{P}\big{\{}a_{1}<p\leq a_{2}|p>a_{1}\big{\}}=$
-$(a_{2}-a_{1})/(1-a_{1})=$
-$a_{2}(1-a_{1}/a_{2})/(1-a_{1})<$
-$a_{2}(1-a_{1})/(1-a_{1})=a_{2}.$
-A.2 Proof of Proposition 2
-Proof.
-By (4), under $H_{0}$ we have
-$r_{\alpha}=\mathbb{E}(\phi_{\alpha}/\alpha)=$
-$\mathbb{P}(p\leq a_{1})\mathbb{E}(\phi_{\alpha}/\alpha|p\leq a_{1})+$
-$\mathbb{P}(a_{1}<p\leq a_{2})\mathbb{E}(\phi_{\alpha}/\alpha|a_{1}<p\leq a_{2})=$
-$a_{1}(1/a_{1})+(a_{2}-a_{1})(1/a_{2})=$
-$1+(a_{2}-a_{1})/a_{2}.$
-Appendix B Multiple hypotheses
-Here we discuss scenarios where $\alpha$ is chosen post hoc and multiple hypotheses are tested.
-Consider the situation where a particular researcher tests multiple hypotheses $H_{1},...,H_{m}$ and computes corresponding p-values $p_{1},...,p_{m}$.
-To avoid a flood of false positive findings, it is common in this setting to apply a multiple testing corrrection.
-Here we will consider control of the familywise error rate (FWER) by use of the Bonferroni method.
-The Bonferroni method takes the raw, i.e.
-unadjusted, p-values $p_{1},...,p_{m}$ as input and outputs adjusted p-values, which are simply $p_{1}^{b}=\min\{p_{1}\cdot m,1\},...,p_{m}^{b}=\min\{p_{m}\cdot m,1\}$.
-The method then rejects the hypotheses with adjusted p-values at most $\alpha$.
-Of course, an equivalent way to see which hypotheses are rejected is to simply check which of the raw p-values are at most $\alpha/m$.
-The number of false positives, i.e.
-the number of true hypotheses that are rejected, if often denoted by $V$.
-The FWER is defined as $\mathbb{P}(V>0)$ and controlling the FWER means ensuring that $\mathbb{P}(V>0)\leq\alpha$.
-The Bonferroni method controls the FWER regardless of the dependence structure of the $m$ p-values and regardless of how many hypotheses are in fact false.
-It should furter be remarked that the Bonferroni procedure can be made slightly more powerful through a stepwise procedure by Holm (Holm, 1979; Goeman and Solari, 2014).
-This method checks which hypotheses are rejected by Bonferroni and removes these from consideration.
-Then Bonferroni is applied to the hypotheses that are left, which can lead to additional rejections and this is repeated until a step is reached where there are no additional rejections.
-Here we will simply consider Bonferroni without the Holm improvement, since if Bonferroni does not control the FWER then by implication neither does Holm.
-While there are many scenarios with data-dependent $\alpha$ where Bonferroni fails miserably, here we will focus on two scenarios, which bear slight resemblance to the scenarios considered in §2.
-As is usual, we will assume that the raw p-values corresponding to the true hypotheses are uniformly distributed on $[0,1]$.
-Further, for the sake of the below examples, we assume that $p_{1},...,p_{m}$ are independent of each other.
-Note that $FWER\leq\alpha$ is equivalent to $\mathbb{E}(\mathbbm{1}_{E})\leq\alpha$, where $E$ is defined to be the event that $V>0$.
-Further note that $\mathbb{E}(\mathbbm{1}_{E})\leq\alpha$ is equivalent to $\mathbb{E}(\mathbbm{1}_{E}/\alpha)\leq 1$, which we use below.
-B.1 Example 1: two possible significance levels $\alpha$
-As in the main paper, suppose the researcher chooses from two thresholds $10<a_{1}<a_{2}<$ 1.
-Let $p_{(1)}=\min\{p_{i}:1\leq i\leq m\}$ denote the smallest raw p-value and let $p_{(1)}^{b}=\min\{mp_{(1)},1\}$ be the smallest Bonferroni-adjusted p-value.
-Similar to before, one can image that a researcher is more likely to use the threshold $a_{1}$ if there is at least one rejection at $\alpha=a_{1}$, i.e.
-if $p_{(1)}^{b}\leq a_{1}$.
-To simplify, suppose she chooses $\alpha=a_{1}$ if $p_{(1)}^{b}\leq a_{1}$ and $\alpha=a_{2}$ otherwise, i.e.
-if $a_{1}<p_{(1)}^{b}\leq 1$.
-Like before, $\alpha$ depends on the p-value and thus the data, but we write $\alpha$ instead of $\alpha(X)$ for brevity.
-B.1.1 No type I error control conditional on $\alpha$
-Suppose all $m$ hypotheses are true and recall that $E=\{V>0\}$ is the event that there is at least one false positive finding.
-Then if $\alpha=a_{1}$, we have $p_{(1)}^{b}\leq a_{1}$.
-Thus $\mathbb{P}(E|\alpha=a_{1})$ is 1, i.e., the FWER is not controlled at all conditional on $\alpha=a_{1}$.
-The following proposition states that conditional on $\alpha=a_{2}$ it is the other way around, i.e.
-the test is conservative: $\mathbb{P}(E|\alpha=a_{2})<a_{2}$.
-Proposition 3.
-Suppose $m\geq 2$ and all hypotheses $H_{1},...,H_{m}$ are true. For the above procedure that takes $\alpha=a_{1}$ if $p\leq a_{1}$ and $\alpha=a_{2}$ otherwise, we have
-$\mathbb{P}(E|\alpha=a_{2})=1-\Big{\{}(1-a_{2}/m)^{m}\big{/}(1-a_{1}/m)^{m}\Big{\}}<a_{2}.$
-Proof.
-Part 1. We have
-$\mathbb{P}(E|\alpha=a_{2})=\mathbb{P}(p_{(1)}^{b}\leq a_{2}|p_{(1)}^{b}>a_{1})=\mathbb{P}(a_{1}<p_{(1)}^{b}\leq a_{2}|p_{(1)}^{b}>a_{1})=$
-$\mathbb{P}(a_{1}<mp_{(1)}\leq a_{2}|mp_{(1)}>a_{1})=\mathbb{P}(a_{1}/m<p_{(1)}\leq a_{2}/m|p_{(1)}>a_{1}/m).$ (7)
-If $m$ variables are independent and have a standard uniform distribution, the minimum of the $m$ variables has a $Beta(1,m)$ distribution (Hemerik et al., 2019), which has density function $f(x)=\{1/B(1,m)\}(1-x)^{m-1}$, where $B(1,m)$ is the normalization constant
-$\int_{0}^{1}(1-x)^{m-1}dx=[-(1-x)^{m}/m]\Big{|}_{0}^{1}=0--1/m=1/m.$
-Thus, $p_{(1)}$ has a $Beta(1,m)$ distribution, so that (7) equals
-[…]
-$\Big{\{}-(1-x)^{m}/m\Big{|}_{a_{1}/m}^{a_{2}/m}\Big{\}}\Big{/}\Big{\{}-(1-x)^{m}/m\Big{|}_{a_{1}/m}^{1}\Big{\}}=$
-$\Big{\{}-(1-a_{2}/m)^{m}--(1-a_{1}/m)^{m}\Big{\}}\Big{/}\Big{\{}0--(1-a_{1}/m)^{m}\Big{\}}=$
-$\Big{\{}(1-a_{1}/m)^{m}-(1-a_{2}/m)^{m}\Big{\}}\Big{/}\Big{\{}(1-a_{1}/m)^{m}\Big{\}}=$
-$1-\Big{\{}(1-a_{2}/m)^{m}\big{/}(1-a_{1}/m)^{m}\Big{\}},$
-as was to be shown.
-Part 2. We now need to show that the above is smaller than $a_{2}$. This is equivalent to showing that
-$(1-a_{2}/m)^{m}/(1-a_{1}/m)^{m}<1-a_{2},$
-which is equivalent to
-$(m-a_{2})^{m}(1-a_{2})^{-1}(m-a_{1})^{-m}<1.$
-The left side is $1$ if $a_{2}=a_{1}$ and we need to show that it is strictly larger than 1 if $a_{2}>a_{1}$. We will do this by showing that the derivative of the function
-$g(a_{2})=(m-a_{2})^{m}(1-a_{2})^{-1}(m-a_{1})^{-m}$
-is strictly positive on $(0,1)$. We use the product rule and the chain rule to compute the derivative:
-$\frac{d}{da_{2}}g(a_{2})=\Big{\{}(m-a_{2})^{m}(--(1-a_{2})^{-2})-m(m-a_{2})^{m-1}(1-a_{2})^{-1}\Big{\}}(m-a_{1})^{-m}.$
-Thus, to show that $\frac{d}{da_{2}}g(a_{2})>0$, it suffices to show that
-$(m-a_{2})^{m}(1-a_{2})^{-2}>m(m-a_{2})^{m-1}(1-a_{2})^{-1},$
-i.e., that
-$(m-a_{2})(1-a_{2})^{-1}>m.$
-This is true since
-$(1-a_{2}/m)(1-a_{2})^{-1}>1,$
-so that we are done.
-So far, we have seen that $\mathbb{P}(E|\alpha=a_{2})<a_{2}$ and $\mathbb{P}(E|\alpha=a_{1})=1>a_{1}$.
-Thus, conditional on $\alpha=a_{2}$ the method is conservative and conditional on $\alpha=a_{1}$ the method is much too liberal.
-Thus, the method does not provide FWER control conditional on the data-dependent $\alpha$.
-B.1.2 No type I error control in terms of expected discrepancy
-The above raises the question question whether we at least control the FWER on average.
-As in the main paper, we can consider the expected value of the discrepancy ratio, which in the context of the FWER is
-$\mathbb{E}(r_{\alpha}):=\mathbb{E}\big{\{}\mathbb{E}(\mathbbm{1}_{E})/\alpha|\alpha\big{\}}=\mathbb{E}(\mathbbm{1}_{E}/\alpha).$
-If this value is at most 1, then we control the FWER in an average sense.
-However, the following proposition implies that $\mathbb{E}(\mathbbm{1}_{E}/\alpha)$ is often larger than 1.
-The reason that it is not always larger than 1 is due to the fact that Bonferroni is conservative under independence if $\alpha$ is fixed, i.e., when we take $a_{1}=a_{2}$.
-Proposition 4.
-In the setting of Proposition 3, we have $\mathbb{E}(\mathbbm{1}_{E}/\alpha)=$
-$\Big{\{}1-(1-a_{2}/m)^{m}\Big{\}}a_{2}^{-1}+\Big{\{}1-(1-a_{1}/m)^{m}\Big{\}}(a_{1}^{-1}-a_{2}^{-1}),$ (8)
-which can be larger than $1$, depending on $m$, $a_{1}$ and $a_{2}$.
-The first term in (8) is smaller than 1 and represents $\mathbb{E}(\mathbbm{1}_{E}/\alpha)$ in the situation that $a_{1}$ and $a_{2}$ are identical, i.e.
-the case that $\alpha$ is fixed.
-The second term represents the increase of $\mathbb{E}(\mathbbm{1}_{E}/\alpha)$ due to $a_{2}$ being larger than $a_{1}$.
-Proof.
-We have
-$\mathbb{E}(\mathbbm{1}_{E}/\alpha)=\mathbb{P}\Big{(}p_{(1)}^{b}\leq a_{1}\Big{)}(1/a_{1})+\mathbb{P}\Big{(}a_{1}\leq p_{(1)}^{b}\leq a_{2}\Big{)}(1/a_{2})=$
-$\mathbb{P}\Big{(}p_{(1)}\leq a_{1}/m\Big{)}(1/a_{1})+\mathbb{P}\Big{(}a_{1}/m\leq p_{(1)}\leq a_{2}/m\Big{)}(1/a_{2})=$
-Like in the proof of Proposition 3, $p_{(1)}$ has a $Beta(1,m)$ distribution, so that the above equals
-[…]
-Recall from the proof of Proposition 3 that $B(1,m)=1/m$. Consequently, the above equals
-$ma_{1}^{-1}\big{\{}-(1-x)^{m}/m\big{\}}\Big{|}_{0}^{a_{1}/m}+ma_{2}^{-1}\big{\{}-(1-x)^{m}/m\big{\}}\Big{|}_{a_{1}/m}^{a_{2}/m}=$
-$ma_{1}^{-1}\Big{\{}-(1-a_{1}/m)^{m}/m--(1-0)^{m}/m\Big{\}}+ma_{2}^{-1}\Big{\{}-(1-a_{2}/m)^{m}/m--(1-a_{1}/m)^{m}/m\Big{\}}=$
-$a_{1}^{-1}\Big{\{}1-(1-a_{1}/m)^{m}\Big{\}}+a_{2}^{-1}\Big{\{}(1-a_{1}/m)^{m}-(1-a_{2}/m)^{m}\Big{\}}$
-$a_{1}^{-1}\Big{\{}1-(1-a_{1}/m)^{m}\Big{\}}+a_{2}^{-1}\Big{\{}(1-(1-a_{2}/m)^{m})-(1-(1-a_{1}/m)^{m})\Big{\}}=$
-$\Big{\{}1-(1-a_{1}/m)^{m}\Big{\}}(a_{1}^{-1}-a_{2}^{-1})+\Big{\{}1-(1-a_{2}/m)^{m}\Big{\}}a_{2}^{-1}.$ (9)
-If we plug in $m=100$, $a_{1}=0.8$ and $a_{2}=0.81$, then the above is 0.70, rounded off to two decimals.
-This can be understood by noting that if we take $a_{1}=a_{2}=0.8$, we actually get a similar number, since we then consider standard Bonferroni with $\alpha=0.8$ and that method is particularly conservative under independence when $\alpha$ is large.
-If we plug in $a_{1}=0.05$ and $a_{2}=0.1$, then (9) is 1.44, rounded off to two decimals. This illustrates that for many values of $a_{1}$ and $a_{2}$, (9) exceeds 1. ∎
-For most reasonable settings, we find with Proposition 4 that $\mathbb{E}(\mathbbm{1}_{E}/\alpha)>1$.
-For example, if $m=1000$, $a_{1}=0.005$ and $a_{2}=0.05$ we find $\mathbb{E}(\mathbbm{1}_{E}/\alpha)\approx 1.87.$ Thus, usually Bonferroni will too liberal, in an average sense.
-B.2 Example 2: infinitely many significance levels $\alpha$
-Here, we again consider $0<C\leq 1$ and assume that if $p_{(1)}^{b}\leq C$, the researcher chooses $\alpha=p_{(1)}^{b}$.
-If $p_{(1)}^{b}>C$, she takes $\alpha=C$.
-Suppose that all $m$ hypotheses are true.
-Then, with this definition of $\alpha$, there is always at least one incorrect rejection if $p_{(1)}^{b}\leq C$.
-Thus, conditional on $\alpha<C$, $\mathbbm{1}_{E}$ is always equal to 1.
-Hence, we clearly do not have $\mathbb{E}(\mathbbm{1}_{E}/\alpha|\alpha<C)\leq 1$, i.e., we do not have FWER control conditional on $\alpha<C$.
-Next we can investigate whether we at least control the FWER on average in the sense that $\mathbb{E}(r_{\alpha})=\mathbb{E}(\mathbbm{1}_{E}/\alpha)\leq 1.$ The following proposition states that this is not the case.
-In fact, $\mathbb{E}(\mathbbm{1}_{E}/\alpha)$ is infinity, like we saw in the case of one hypothesis.
-Proposition 5.
-Consider the setting of Propositions 3 and 4 above, but define $\alpha=p_{(1)}^{b}$ if $p_{(1)}^{b}\leq C$ and $\alpha=C$ otherwise. Then
-$\mathbb{E}(r_{\alpha})=\mathbb{E}(\mathbbm{1}_{E}/\alpha)=\infty.$
-Proof.
-We will again use that $p_{(1)}$ has a $Beta(1,m)$ distribution, so has density function $f(x)=B(1,m)(1-x)^{m-1}$.
-Note that $E$ happens if and only if $p_{(1)}^{b}\leq C$, i.e., if and only if $p_{(1)}\leq C/m$.
-Hence, $\mathbb{E}(\mathbbm{1}_{E}/\alpha)=$
-[…]
-where we used that if $p_{(1)}^{b}\leq C$ then $1/\alpha=1/(mp_{(1)})$. Since $B(1,m)^{-1}=m$, the above equals
-[…]
-Note that $(1-x)^{m-1}\geq(1-C/m)^{m-1}$ for all $0\leq x\leq C/m$. Hence the above is at least
-[…]